# Unsupervised learning in particle physics

### v1

@@ Appendix A Kullback-Liebler divergence
Kullback-Liebler ($\mathbb{KL}$) divergence is a measure of how a probability distribution diverges from an expected probability distribution. In simple words, we can think it of as a measure that shows how different the two distributions are from each other. The $\mathbb{KL}$ divergence can be used for both discrete and continuous distributions. Given two discrete probability distributions, $\rho$ (actual distribution) and $\pi$ (reference distribution), the $\mathbb{KL}$ divergence is calculated as the sum over all possible outcomes:
@@ Appendix B Jensen-Shannon divergence
The Jensen-Shannon divergence is another measure of similarity between two probability distributions. In the context of GANs, the JSD is often used to quantify the difference between the distribution of real data and the distribution of generated data. Let $\rho$ represent the distribution of real data and $\pi$ represent the distribution of generated data from a GAN. The Jenson-Shannon Divergence between $\rho$ and $\pi$ is defined as:

### v3

@@ Appendix A Kullback-Liebler divergence
Kullback-Liebler ($\mathbb{KL}$) divergence [150] is a measure of how a probability distribution diverges from an expected probability distribution. In simple words, we can think it of as a measure that shows how different the two distributions are from each other. The $\mathbb{KL}$ divergence can be used for both discrete and continuous distributions. Given two discrete probability distributions, $\rho$ (actual distribution) and $\pi$ (reference distribution), the $\mathbb{KL}$ divergence is calculated as the sum over all possible outcomes:
@@ Appendix B Jensen-Shannon divergence
The Jensen-Shannon divergence [151, 152] is another measure of similarity between two probability distributions. In the context of GANs, the JSD is often used to quantify the difference between the distribution of real data and the distribution of generated data. Let $\rho$ represent the distribution of real data and $\pi$ represent the distribution of generated data from a GAN. The Jenson-Shannon Divergence between $\rho$ and $\pi$ is defined as: